OpenStudy (anonymous):

How can you tell when a quadratic equation has two identical, rational solutions? when the radicand is negative when b in the quadratic formula is greater than the radicand when the radicand equals zero when the radicand is not a perfect square

OpenStudy (anonymous):

If the radicand is 0 then the solution would like something like this: \[x = -b +/- \sqrt{0}\] -b + 0 = -b - 0 Which means that the roots are identical when the radicand is zero

OpenStudy (anonymous):

so the 3rd one

OpenStudy (anonymous):

Yes, thats correct!